40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 25 02:03:59 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts 
    Reply  Fri Jul 28 15:36:32 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts errSig.pdfQPD_simulink.pngfeedbackTopology.pdf
       Reply  Tue Aug 1 16:05:01 2017, gautam, Update, Optical Levers, Optical lever tuning - cost function construction loopPerformance.pdfOplevLoop.pngOL_errSigs.pdfDAC_saturation.pdf
          Reply  Thu Nov 16 13:57:01 2017, gautam, Update, Optical Levers, Optical lever noise VertexOLnoise.pdfITMYvsSRM.pdf
             Reply  Thu Nov 16 15:43:01 2017, rana, Update, Optical Levers, Optical lever noise 
                Reply  Tue Nov 21 11:37:29 2017, gautam, Update, Optical Levers, BS OL calibration updated OL_calib_BS_PERROR.pdfOL_calib_BS_YERROR.pdfVertexOLnoise_updated.pdf
                   Reply  Tue Nov 21 16:28:23 2017, gautam, Update, Optical Levers, BS OL calibration updated BS_oplev_sweep.png
                      Reply  Tue Nov 21 23:04:12 2017, gautam, Update, Optical Levers, Oplev "noise budget" vertexOL_noises.pdf
                         Reply  Wed Nov 22 05:41:32 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                            Reply  Wed Nov 22 15:29:23 2017, gautam, Update, Optical Levers, Oplev "noise budget" OplevTest.jpg
                               Reply  Wed Nov 22 16:40:00 2017, Koji, Update, Optical Levers, Oplev "noise budget" 
                               Reply  Wed Nov 22 19:20:01 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                                  Reply  Wed Nov 22 23:56:14 2017, gautam, Update, Optical Levers, Oplev "noise budget" 
                                     Reply  Thu Nov 23 18:03:52 2017, gautam, Update, Optical Levers, Oplev "noise budget" ITMY_P_noise.pdfITMY_Y_noise.pdf
Message ID: 13432     Entry time: Thu Nov 16 13:57:01 2017     In reply to: 13156     Reply to this: 13433
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: Optical lever noise 

I disabled the OL loops for ITMX, ITMY and BS at GPStime 1194897655 to come up with an Oplev noise budget. OL spots were reasonably well centered - by that, I mean that the PIT/YAW error signals were less than 20urad in absolute value.

Attachment #1 is a first look at the DTT spectra - I wonder why the BS Oplev signals don't agree with the ITMs at ~1Hz? Perhaps the calibration factor is off? The sensing noise not really flat above 100Hz - I wonder what all those peaky features are. Recall that the ITM OLs have analog whitening filters before the ADC, but the BS doesn't...

In Attachment #2, I show comparison of the error signal spectra for ITMY and SRM - they're on the same stack, but the SRM channels don't have analog de-whitening before the ADC.

For some reason, DTT won't let me save plots with latex in the axes labels...

Attachment 1: VertexOLnoise.pdf  78 kB  | Hide | Hide all
VertexOLnoise.pdf
Attachment 2: ITMYvsSRM.pdf  51 kB  | Hide | Hide all
ITMYvsSRM.pdf
ELOG V3.1.3-